OpenStudy (anonymous):

What is the value of Dx in the following system? x + 3y = –3 2x + 5y = –7 A. -36 B. -1 C.-13 D.6

OpenStudy (anonymous):

@johnweldon1993 can you help?

OpenStudy (johnweldon1993):

The value of "Dx"? Are you sure you don't mean just 'x' ?

OpenStudy (anonymous):

D sub X

OpenStudy (johnweldon1993):

Sorry just reviewing discriminants... hang on :)

OpenStudy (johnweldon1993):

Okay..that's right too...to find the discriminant of a system of linear equations x + 3y = –3 2x + 5y = –7 Dx= (-3 * 5) - (3 * -7) Dx = -15 - (-21) Dx = -15 + 21 Dx = 6 Let me think about how to explain it though ...:)

OpenStudy (anonymous):

Okay I understand! thanks for making it easy to follow. :)